# Privacy Guarantees in Posterior Sampling under Contamination

### v2

@@ C.4 Quantile Regression Case
$\succeq$ $\,\|\theta-\theta^{*}\|_{2}^{2}$
where $\tau^{*}=\max\{\tau,1-\tau\}$, $a_{3}=\max\{1-a_{1},1-a_{2}\}$. The last step is by applying the covariance SVD argument as in the end of Lemma 22. ∎
Appendix D Additional Discussions
D.1 For Strictly Adjacent Dataset
@@
Two datasets $\mathcal{D}$ and $\bm{X}$ are strictly adjacent if and only if $\exists\bm{x}\notin\mathcal{D}$ such that $\bm{X}=\mathcal{D}\cup\{\bm{x}\}$.
@@
Another definition for neighbouring datasets exists in the literature where one is a subset of another, and the larger dataset contains exactly one more entry.
Using the notation above, $\bm{X}=\mathcal{D}\cup\{\bm{x}\}$, so $\mathcal{D}$ and $\bm{X}$ is a pair of neighbouring datasets. In this case, (4) becomes either
$\mathbb{P}_{\pi_{n}}(S|\bm{X})\leq\sqrt{\eta_{n}}\left[\sqrt{\eta_{n}}+\frac{m_{n-1}(A_{n}^{c},\mathcal{D})}{m_{n}(\Omega,\bm{X})}\right]\mathbb{P}_{\pi_{n-1}}(S|\mathcal{D})+\frac{m_{n-1}(A_{n}^{c},\mathcal{D})}{m_{n}(\Omega,\bm{X})}.$
@@
$\mathbb{P}_{\pi_{n-1}}(S|\mathcal{D})\leq\sqrt{\eta_{n}}\left[\sqrt{\eta_{n}}+\frac{m_{n}(A_{n}^{c},\bm{X})}{m_{n-1}(\Omega,\mathcal{D})}\right]\mathbb{P}_{\pi_{n}}(S|\bm{X})+\frac{m_{n}(A_{n}^{c},\bm{X})}{m_{n-1}(\Omega,\mathcal{D})}.$
@@
Under Assumptions 1-5 for any pair of strictly adjacent datasets $\bm{X}$ and $\bm{Z}$, regardless of which one is a larger set. With slight abuse of notation, let $\pi_{*}$ denote the appropriate posterior distribution $\pi_{n}$ or $\pi_{n-1}$ depending on the size of the given dataset, we have
@@
where $\epsilon_{n}$ is defined with respect to $\tfrac{1}{2}\varepsilon_{n}$ and $\delta_{n}$, while $\varepsilon_{n}$,$\delta_{n}$, $t_{n}$, $c_{5}$ are defined as in Theorem 5.
@@
Let $n=\max\{|\bm{X}|,|\bm{Z}|\}$. Recall that the term $m_{n}(\Omega,\cdot)$ only depend on the sequence $t_{n}$ and $m_{n}(A_{n}^{c},\cdot)$ only depend on $\phi_{n}$. Denote the lagged sequences $Lt_{n}=t_{n-1}$, $L\phi_{n}=\phi_{n-1}$, and the scaled sequences $\tilde{t}_{n}=\tfrac{n}{n-1}t_{n}$, $\tilde{\phi}_{n}=\sqrt{\tfrac{n}{n-1}}\phi_{n}$. For instance, if we want to evaluate $\tfrac{m_{n-1}(A_{n}^{c},\mathcal{D})}{m_{n}(\Omega,\bm{X})}$, we can evaluate the $m_{n}(\Omega,\bm{X})$ on $Lt_{n}$ but $m_{n-1}(A_{n}^{c},\mathcal{D})$ on $\tilde{\phi}_{n-1}$, then the indices in the obtained result are synchronized to $n-1$.
Since only one sequence needs to be lagged, without loss of generality, we can always synchronize $n-1$ and thus the rest follows. ∎
D.2 Privacy in Non-identifiable Models
@@
The ”Identifiability Assumption” does not change the level of differential privacy measured in ($\epsilon$,$\delta$). However, to determine the convergence rate for ($\epsilon$,$\delta$) in Theorem 5 requires sufficient knowledge of the model, including how the density functions are indexed.
In other words, it is not crucial to have an identifiable parameterization, in order for the model to be differentially private. In fact, under suitable regularity conditions, the level of privacy in terms of $\epsilon$ or ($\epsilon,\delta$) is intrinsic to the model, regardless of how its parameter space and likelihood function take form. The level of privacy is preserved by applying a suitable transformation (not necessarily bijective) to the parameter space and the prior. Hence, the resulting model can be considered as a reparameterization of the original model.
@@
Let $(\Omega_{\psi},\Sigma_{\psi},\mathtt{d}\psi)$ and $(\Omega_{\theta},\Sigma_{\theta},\mathtt{d}\theta)$ be two measure spaces and a diffeomorphism $t:\Omega_{\psi}\rightarrow\Omega_{\theta}$ such that $t^{-1}$ is well-defined and differentiable. Then the differential privacy level of model $f_{\psi}(\bm{x};\psi)$ parameterized by $\Omega_{\psi}$ with prior $\pi_{0}(\psi)$ is the same as the privacy level of model $f_{\theta}(\bm{x};\theta)$ parameterized by $\Omega_{\theta}$ with prior $\tilde{\pi}_{0}(\theta)$,
@@
Recall that $f_{\psi}(\bm{x};\psi)$ and $\pi_{0}(\psi)$ are the likelihood and the prior with respect to the parameter space $\Omega_{\psi}$, and,
where $J_{t^{-1}}(\theta)$ is the Jacobian of the inverse function. Note that $f_{\theta}$ and $\tilde{\pi}_{0}$ are derived from $f_{\psi}$, $\pi_{0}$ through change of measure under mapping $t$.
@@
$\implies$ $\,\mathbb{P}_{\pi_{\psi}}(t^{-1}(S)|\bm{X})\leq e^{\epsilon}\mathbb{P}_{\pi_{\psi}}(t^{-1}(S)|\bm{Z})+\delta$
and vice versa, since the integral is preserved through the change of measure. Therefore, a pair of $(\epsilon,\delta)$ holds for a model parameterized in $\Omega_{\psi}$ if and only if it holds for that model parameterized in $\Omega_{\theta}$ through the diffeomorphism $t$. ∎
@@
where $\mu$ is the base measure for the measure space $\Omega_{\psi}$. That is $\psi_{1}$ and $\psi_{2}$ lead to two likelihood functions that are equal almost everywhere. Since equal a.e. could create an issue in the DP inequality since the worst case could land on the discontinuity, we consider $f(\bm{x};\psi)$ to be continuous in $\bm{x}$. In which case, $\psi_{1}\sim\psi_{2}$ if and only if they represent the same likelihood function on the whole of $\mathcal{X}$.
@@
$\begin{array}[]{cccc}t:&\Omega_{\psi}&\rightarrow&\Omega_{\theta}\\ &\psi&\mapsto&[\psi]\end{array}$
and if there is a way to choose every representative $[\psi]$ such that $t$ is measurable, then we have the following proposition.
@@
Let $(\Omega_{\psi},\Sigma_{\psi},\mathtt{d}\psi)$ be a measure space such that the model with likelihood function $f_{\psi}(x;\psi)$ and prior $\pi_{0}(\psi)$ is non-identifiable. Suppose that the quotient map $t:\Omega_{\psi}\rightarrow\Omega_{\theta}\subset\Omega_{\psi}$ that reparameterizes the model onto the measure space $(\Omega_{\theta},\Sigma_{\theta},\mathtt{d}\theta)$ is measurable, then the model parameterized in $\Omega_{\theta}$ has the same level of differential privacy as the model parameterized in $\Omega_{\psi}$.
@@
under the same construction of the quotient map $t$, pushforward measure $t_{*}(\mathbb{P}_{\pi_{0}})$ and its density $\tilde{\pi}_{0}$ with respect to $\mathtt{d}\theta$ as in the previous proof. Since the pushforward measure and the design of the quotient ensure any integrals are preserved, the $\implies$ direction is trivial.
@@
$\frac{\pi_{\psi}(\psi|\bm{X})}{\pi_{\psi}(\psi|\bm{Z})}<e^{\epsilon},\quad\forall\psi\in B$
@@
Without loss of generality, it is now sufficient to check that $(\epsilon,\delta)$ hold for all sets $A\in\Sigma_{\psi}$ such that
@@
For any set $A\in\Sigma_{\psi}$ such that (38) holds. Let $S=t(A)$, note that any $\psi\in t^{-1}(S)\supseteq S$ also satisfies (38),
@@
{funding}
All authors were supported by the EPSRC research grant ”Pooling INference and COmbining Distributions Exactly: A Bayesian approach (PINCODE)”, reference (EP/X028100/1, EP/X028119/1, EP/X028712/1, EP/X027872/1).
LA, HD, MP and GOR were also supported by by the UKRI grant, ”On intelligenCE And Networks (OCEAN)”, reference (EP/Y014650/1).

### v3

@@ C.4 Quantile Regression Case
$\succeq$ $\,\|\theta-\theta^{*}\|_{2}^{2},$
where $\tau^{*}=\max\{\tau,1-\tau\}$, $a_{3}=\max\{1-a_{1},1-a_{2}\}$. The last step is to apply the covariance SVD argument as in the end of Lemma 21. ∎
Appendix D
D.1 Alternative Proof for Posterior Contraction
The proof strategy for our main result is roughly divided into two parts:
  1. 1.
    Bounding the likelihood ratio within a compact neighbourhood around $\theta^{*}$;
  2. 2.
    Bounding the posterior mass of the set complement of that neighbourhood.
The second component of our analysis involves establishing the contraction rate of the posterior distribution, following the work of Wong and Shen (1995); Shen and Wasserman (2001). Posterior contraction rates were also investigated independently in Ghosal et al. (2000), and Ghosal and Van der Vaart (2017) presents results that extend beyond the i.i.d. observation setting. In what follows, we briefly discuss the assumptions and conclusions of our work when the proof strategy of Theorem 8.11 in Ghosal and Van der Vaart (2017) is adopted instead. Please note that all the constants in this section are independent of the constants in the main text, even when the same symbols are used, although some are functionally equivalent.
Definition 12 (Test).
Let $(\mathbb{X},\mathcal{X})$ be a measurable space. A test function $\varphi$ defined on $(\mathbb{X},\mathcal{X})$ is a measurable function $\varphi:\mathbb{X}\rightarrow[0,1]$.
Let $h^{2}(\theta,\theta^{\prime})$ denote the squared Hellinger distance between two likelihood functions indexed by $\theta$ and $\theta^{\prime}$, i.e., $f(\cdot;\theta)$ and $f(\cdot;\theta^{\prime})$. We are not concerned with the contaminated likelihood $k_{p}$ since the assumptions to be stated below all hold for any $p>0$ whenever they hold for $p=0$, i.e., for $f$.
Assumption 6 (Basic Testing Assumption).
Let $\xi,K>0$ be some universal constants. For every $n\in\mathbb{N}$, $\phi>0$ and $\theta$ with $h(\theta,\theta^{*})>\phi$, there exists a test $\varphi_{n}:\mathbb{X}^{n}\rightarrow[0,1]$ such that
$\mathbb{E}_{\theta^{*}}[\varphi_{n}]\leq\exp(-Kn\phi^{2}),\qquad\sup_{h(\theta^{\prime},\theta)\leq\xi\phi}\mathbb{E}_{\theta^{\prime}}[1-\varphi_{n}]\leq\exp(-Kn\phi^{2}).$ (35)
By Proposition D.8 of (Ghosal and Van der Vaart, 2017), the above assumption is always satisfied for any likelihood model with universal constants $\xi=1/2$, $K=1/8$.
Recall that $\Omega$ denotes the parameter space of interest. Let $B_{KL}(\theta,\phi):=\{\theta^{\prime}\in\Omega:KL(\theta\|\theta^{\prime})<\phi\}$ denote the $\phi$-neighbourhood of $\theta$ in KL-divergence.
Assumption 7.
There exists a partition of $\Omega$, namely $\Omega_{n}\cup\Omega_{n}^{c}$ such that for constants $\phi_{n},\bar{\phi}_{n}\geq n^{-1/2}$, and every sufficiently large $j$,
  1. (i)
    $\frac{\mathbb{P}_{\pi_{0}}(\theta:j\phi_{n}<h(\theta,\theta^{*})\leq 2j\phi_{n})}{\mathbb{P}_{\pi_{0}}(B_{KL}(\theta^{*},\phi_{n}))}\leq e^{Kj^{2}n\phi_{n}/2},$
  2. (ii)
    $\sup_{\phi\geq\phi_{n}}\log N(\xi\phi,\{f(\cdot;\theta):\theta\in\Omega_{n},h(\theta,\theta^{*})\leq 2\phi\})\leq n\phi_{n}^{2},$
    where $N$ denotes the covering number under the Hellinger distance.
  3. (iii)
    There exists some constant $C>0$ such that
    $\frac{\mathbb{P}_{\pi_{0}}(\Omega_{n}^{c})}{\mathbb{P}_{\pi_{0}}(B_{KL}(\theta^{*},\bar{\phi}_{n}))}=o\left(e^{-D_{n}n\bar{\phi}_{n}^{2}}\right),$
    for some $D_{n}\rightarrow\infty$.
Define the neighbourhood set $A_{n}\subset\Omega$ by $A_{n}(r):=\{\theta\in\Omega:h(\theta,\theta^{*})\geq r\phi_{n}\}$. Note that the size of the neighbourhood $A_{n}$ depends on an additional value $r$.
Theorem 29 (Theorem 8.11 (Ghosal and Van der Vaart, 2017)).
Under Assumptions 6 and 7,
$\mathbb{P}_{\pi_{n}}(A_{n}(M_{n})|\bm{X})\rightarrow 0\quad\text{ in $\mathbb{P}_{\theta^{*}}^{(n)}$-probability}$
for any arbitrarily slow $M_{n}\rightarrow\infty$.
More specifically, there exists some constant $C_{1}>0$ such that for any $n$ large enough,
$\mathbb{P}_{\theta^{*}}^{(n)}\left(\mathbb{P}_{\pi_{n}}(A_{n}(M_{n})|\bm{X})>e^{-\tfrac{1}{2}(KM_{n}^{2}-1)n\phi_{n}^{2}}+e^{-(\tfrac{1}{4}KM_{n}^{2}-D)n\phi_{n}^{2}}+C_{1}e^{-\tfrac{1}{2}(D_{n}-2D)n\bar{\phi}_{n}^{2}}\right)\\ <\frac{e^{-\tfrac{1}{2}(KM_{n}^{2}-1)n\phi_{n}^{2}}}{1-e^{-KM_{n}^{2}n\phi_{n}^{2}}}+\sum_{j\geq 1}e^{-(\tfrac{1}{4}KM_{n}^{2}(2j^{2}-1)-D)n\phi_{n}^{2}}+e^{-\tfrac{1}{2}(D_{n}-2D)n\bar{\phi}_{n}^{2}}+\frac{2}{D}$ (36)
holds for any $D>0$.
Sketch Proof..
First, by Assumption 6, (ii) of Assumption 7 and Theorem D.5 of (Ghosal and Van der Vaart, 2017), for given $M_{n}>1$, there exist tests $\varphi_{n}$ such that for every $j\in\mathbb{N}$,
$\mathbb{E}_{\theta^{*}}[\varphi_{n}]\leq e^{n\phi_{n}^{2}}\frac{e^{-KM_{n}^{2}n\phi_{n}^{2}}}{1-e^{-KM_{n}^{2}n\phi_{n}^{2}}},\qquad\sup_{\theta\in\Omega_{n}:h(\theta,\theta^{*})>Mj\phi_{n}}\mathbb{E}_{\theta}[1-\varphi_{n}]\leq e^{-KM_{n}^{2}j^{2}n\phi_{n}^{2}}.$ (37)
As in the proof of Lemma 8, we split $\Omega_{n}$ into countably many shells
$\mathcal{S}_{n,j}:=\{\theta\in\Omega_{n}:M_{n}\\ j\phi_{n}<h(\theta,\theta^{*})\leq M_{n}(j+1)\phi_{n}\},$
thus $\cup_{j\geq 1}\mathcal{S}_{n,j}=\Omega_{n}\cap A_{n}(M_{n})$. For notational simplicity, we will abbreviate $A_{n}(M_{n})$ to just $A_{n}$. Define the event $E_{n}\subset\mathbb{X}^{n}$ such that
$\int_{\Omega}\prod_{i=1}^{n}\frac{k_{p}(X_{i};\theta)}{k_{p}(X_{i};\theta^{*})}\pi_{0}(\theta)\mathtt{d}\theta\geq e^{-2Dn\phi_{n}^{2}}\mathbb{P}_{\pi_{0}}(B_{KL}(\theta^{*},\phi_{n})).$
Consider the same decomposition of the posterior mass
$\mathbb{P}_{\pi_{n}}(A_{n}|\bm{X})\leq\mathbb{P}_{\pi_{n}}(A_{n}\cap\Omega_{n}|\bm{X})+\mathbb{P}_{\pi_{n}}(\Omega_{n}^{c}|\bm{X}).$
The first term may be bounded by
$\,\mathbb{P}_{\pi_{n}}(A_{n}\cap\Omega_{n}|\bm{X})\leq\frac{\int_{A_{n}\cap\Omega_{n}}\prod_{i=1}^{n}\frac{k_{p}(X_{i};\theta)}{k_{p}(X_{i};\theta^{*})}\pi_{0}(\theta)\mathtt{d}\theta}{\int_{\Omega}\prod_{i=1}^{n}\frac{k_{p}(X_{i};\theta)}{k_{p}(X_{i};\theta^{*})}\pi_{0}(\theta)\mathtt{d}\theta}$
$\leq$ $\,\mathbb{I}_{E_{n}^{c}}+\mathbb{I}_{E_{n}}\frac{\int_{A_{n}\cap\Omega_{n}}\prod_{i=1}^{n}\frac{k_{p}(X_{i};\theta)}{k_{p}(X_{i};\theta^{*})}\pi_{0}(\theta)\mathtt{d}\theta}{\int_{\Omega}\prod_{i=1}^{n}\frac{k_{p}(X_{i};\theta)}{k_{p}(X_{i};\theta^{*})}\pi_{0}(\theta)\mathtt{d}\theta}(\varphi_{n}+1-\varphi_{n})$
$\leq$ $\,\mathbb{I}_{E_{n}^{c}}+\varphi_{n}+\frac{\int_{A_{n}\cap\Omega_{n}}\prod_{i=1}^{n}\frac{k_{p}(X_{i};\theta)}{k_{p}(X_{i};\theta^{*})}\pi_{0}(\theta)\mathtt{d}\theta(1-\varphi_{n})}{e^{-2Dn\phi_{n}^{2}}B_{KL}(\theta^{*},\phi_{n})}.$
  1. 1.
    By Theorem 6.26 in Ghosal and Van der Vaart (2017), the event $E_{n}^{c}$ holds with probability at most $1/D$ under $\mathbb{P}_{\theta^{*}}^{n}$, contributing to one portion of the $1/D$ in the failure rate.
  2. 2.
    The expectation of $\varphi_{n}$ under $\mathbb{P}^{n}_{\theta^{*}}$ is bounded by applying (37), which may be turned into a probabilistic bound in $\mathbb{P}^{n}_{\theta^{*}}$ by applying Markov’s inequality, contributing to the first term in the inner inequality of (36) and the first term in the failure probability.
  3. 3.
    The integral in the final term may be split into integrals on $\cup_{j\geq 1}\mathcal{S}_{n,j}$. Thus, the expectation of the last term is bounded using (i) in Assumption 7. Applying Markov’s inequality to get the infinite-sum terms in (36).
Similarly, $\mathbb{P}_{\pi_{n}}(\Omega_{n}^{c}|\bm{X})$ may be bounded by
$\mathbb{P}_{\pi_{n}}(\Omega_{n}^{c}|\bm{X})\leq\mathbb{I}_{E_{n}^{c}}+\frac{\mathbb{P}_{\pi_{n}}(\Omega_{n}^{c}|\bm{X})}{e^{-2Dn\bar{\phi}_{n}^{2}}\mathbb{P}_{\pi_{0}}(B_{KL}(\theta^{*},\bar{\phi}_{n}))}.$
  1. 1.
    Again, by Theorem 6.26 in Ghosal and Van der Vaart (2017), the event $E_{n}^{c}$ holds with probability at most $1/D$ under $\mathbb{P}_{\theta^{*}}^{n}$, contributing to the other portion of the $1/D$ in the failure rate.
  2. 2.
    By (iii) of Assumption 7, the second term above is uniformly bounded for any $\bm{X}$, contributing to the third term in the inner inequality and failure rate of (36).
Remark 17 (Practicality of Assumptions).
The assumptions (i-iii) in Assumption 7 replace Assumptions 2,1,3 in the main text respectively.
  • Assumption 7 (i) is implied by the following condition:
    There exists constant a $C_{2}>0$ such that for constants $\phi_{n}\leq n^{-1/2}$,
    $\mathbb{P}_{\pi_{0}}(B_{KL}(\theta^{*},\phi_{n}))\geq e^{-C_{2}n\phi_{n}^{2}},$ (38)
    which is weaker than our Assumption 2 (Shen and Wasserman, 2001).
  • Assumption 7 (ii) may be checked through Lemma 2 with a slightly altered argument to bound the covering number (in Hellinger distance) instead of the bracketing number.
  • By assuming (38), Assumption 7 (iii) may be checked through Lemma 7 if we take the prior $\pi_{0}$ to be standard Gaussian and $\Omega_{n}$ to be an expanding $L_{2}$-ball. The radius of $\Omega_{n}$ needs to grow polynomially with $n$ for $D_{n}-2D>0$, considering that $D$ is also growing with $n$.
Overall, the assumptions are somewhat equivalent to those made in the main text, with (38) potentially easier to verify than our Assumption 2 in certain models.
Remark 18.
Note that for (36) to be meaningful, it is necessary that $KM_{n}^{2}>4D$ and $D_{n}>2D$ for the quantities to contract with $n$ and $M_{n}\phi_{n}\rightarrow 0$ for the set $A_{n}$ to contract. Since $\phi_{n}$ contracts to 0 no faster than $n^{-1/2}$ by definition, $M_{n}$ cannot grow faster than $O(n^{1/2})$. Consequently, $D$ cannot grow faster than $O(n)$, implying that the failure probability associated with the upper bound on the posterior mass can be shown to decay only at a polynomial rate in $n$. To prove the contraction rate in the strong sense (for almost sure convergence), a stronger version of (38) is required, and we refer the reader to Theorem 8.9 (Ghosal and Van der Vaart, 2017) for more details. However, the failure rate still decreases polynomially with $n$ when invoking Theorem 8.9 (Ghosal and Van der Vaart, 2017).
Remark 19.
The minimal $n$ for Theorem 29 to hold depends on the sequence of $M_{n}$ and the prior $\pi_{0}$, which can be significantly smaller than that of Theorem 5.
D.1.1 Potential Extensions
In this paper, our primary focus has been on finite-dimensional models with i.i.d. observations. However, posterior contraction results exist for a broader range of setups, including independent but not identical observations, misspecified models, and nonparametric models (Ghosal and Van der Vaart, 2017).
For our proof strategy to remain applicable in these extended frameworks, it is essential to identify a suitable contamination density $g(\bm{x})$ that has a heavier tail than the likelihood function $f(\bm{x};\theta)$, where $\theta$ is potentially infinite. Our Assumption 4 requires $g(\bm{x})$ to have a heavier tail than $\nabla_{\theta}f(\bm{x};\theta)$ for any $\theta$ to better characterise the contraction rate of $\epsilon_{n}$. However, this might not be appropriate in the infinite-dimensional setting. Instead, it is more suitable to characterise the contraction rate of $d(\bm{x};\theta)=\tfrac{k_{p}(\bm{x};\theta)}{k_{p}(\bm{x};\theta^{*}_{0})}$, where $\theta^{*}_{0}$ denotes the minimiser, within the parameter space $\Omega$, of the KL divergence from $k_{p}(\cdot;\theta)$ to the true data-generating measure $p^{*}$, which coincides with $\theta^{*}$ if the model is correctly specified. In most cases, the assumptions required to establish posterior contraction remain analogous to those employed in the finite-dimensional i.i.d. case, i.e., a lower bound on prior mass of a shrinking neighbourhood $A_{n}$ around $\theta^{*}_{0}$, an upper bound on prior mass of the complement of an expanding sieve $\Omega_{n}$, and an upper bound for the complexity (in terms of covering number or bracketing number) of the space of likelihood functions indexed by the sieve $\Omega_{n}$. For a comprehensive exposition of these results, we refer the reader to Chapter 8 of Ghosal and Van der Vaart (2017).
Appendix E Additional Discussions
E.1 For Strictly Adjacent Dataset
Definition 1 is perhaps the most adopted definition for neighbouring datasets in the differential privacy literature, where the two datasets are explicitly required to have the same cardinality.
In some places, it is alternatively defined as one dataset being a proper subset of the other with strictly one fewer element, for instance, in Dwork (2008). To avoid ambiguity, we call two datasets $\bm{X}$ and $\bm{Z}$ strictly adjacent if $\bm{Z}=\bm{X}\cup\{\bm{z}\}$ and $\bm{z}\notin\bm{X}$ (or the same statement with $\bm{Z}$ and $\bm{X}$ swapped holds). Our result translates easily to the strictly adjacent case.
@@ E.1 For Strictly Adjacent Dataset
Two datasets $\bm{Z}$ and $\bm{X}$ are strictly adjacent if and only if $\exists\bm{x}\notin\bm{Z}$ such that $\bm{X}=\bm{Z}\cup\{\bm{x}\}$.
@@
Using the notation above, $\bm{X}=\bm{Z}\cup\{\bm{x}\}$, so $\bm{Z}$ and $\bm{X}$ are a pair of neighbouring datasets. In this case, (4) becomes either
$\mathbb{P}_{\pi_{n}}(S|\bm{X})\leq\sqrt{\eta_{n}}\left[\sqrt{\eta_{n}}+\frac{m_{n-1}(A_{n}^{c},\bm{Z})}{m_{n}(\Omega,\bm{X})}\right]\mathbb{P}_{\pi_{n-1}}(S|\bm{Z})+\frac{m_{n-1}(A_{n}^{c},\bm{Z})}{m_{n}(\Omega,\bm{X})}.$
@@
$\mathbb{P}_{\pi_{n-1}}(S|\bm{Z})\leq\sqrt{\eta_{n}}\left[\sqrt{\eta_{n}}+\frac{m_{n}(A_{n}^{c},\bm{X})}{m_{n-1}(\Omega,\bm{Z})}\right]\mathbb{P}_{\pi_{n}}(S|\bm{X})+\frac{m_{n}(A_{n}^{c},\bm{X})}{m_{n-1}(\Omega,\bm{Z})}.$
@@
Under Assumptions 1-5, for any pair of strictly adjacent datasets $\bm{X}$ and $\bm{Z}$, regardless of which one is larger. With slight abuse of notation, let $\pi_{*}$ denote the appropriate posterior distribution $\pi_{n}$ or $\pi_{n-1}$ depending on the size of the given dataset, we have
@@
where $\epsilon_{n}$ is defined with respect to $\tfrac{1}{2}\varepsilon_{n}$ and $\delta_{n}$, while $\varepsilon_{n}$$\delta_{n}$, $t_{n}$, $c_{5}$ are defined as in Theorem 5.
@@
Let $n=\max\{|\bm{X}|,|\bm{Z}|\}$. Recall that the term $m_{n}(\Omega,\cdot)$ only depends on the sequence $t_{n}$ and $m_{n}(A_{n}^{c},\cdot)$ only depends on $\phi_{n}$. Denote the lagged sequences $Lt_{n}=t_{n-1}$, $L\phi_{n}=\phi_{n-1}$, and the scaled sequences $\tilde{t}_{n}=\tfrac{n}{n-1}t_{n}$, $\tilde{\phi}_{n}=\sqrt{\tfrac{n}{n-1}}\phi_{n}$. For instance, if we want to evaluate $\tfrac{m_{n-1}(A_{n}^{c},\bm{Z})}{m_{n}(\Omega,\bm{X})}$, we can evaluate the $m_{n}(\Omega,\bm{X})$ on $Lt_{n}$ but $m_{n-1}(A_{n}^{c},\bm{Z})$ on $\tilde{\phi}_{n-1}$, then the indices in the obtained result are synchronized to n1n-1.
Since only one sequence needs to be lagged, without loss of generality, we can always synchronise to n1n-1 and thus the rest follows. ∎
E.2 Privacy in Non-identifiable Models
@@
The ”Identifiability Assumption” does not change the level of differential privacy measured in ($\epsilon$$\delta$). However, to determine the convergence rate for ($\epsilon$$\delta$) in Theorem 5 requires sufficient knowledge of the model, including how the density functions are indexed.
In other words, it is not crucial to have an identifiable parameterisation in order for the model to be differentially private. In fact, under suitable regularity conditions, the level of privacy in terms of $\epsilon$ or ($\epsilon,\delta$) is intrinsic to the model, regardless of how its parameter space and likelihood function take form. The level of privacy is preserved by applying a suitable transformation (not necessarily bijective) to the parameter space and the prior. Hence, the resulting model can be considered as a reparameterisation of the original model.
@@
Let $(\Omega_{\psi},\Sigma_{\psi},\mathtt{d}\psi)$ and $(\Omega_{\theta},\Sigma_{\theta},\mathtt{d}\theta)$ be two measure spaces and a diffeomorphism $t:\Omega_{\psi}\rightarrow\Omega_{\theta}$ such that $t^{-1}$ is well-defined and differentiable. Then the differential privacy level of model $f_{\psi}(\bm{x};\psi)$ parameterised by $\Omega_{\psi}$ with prior $\pi_{0}(\psi)$ is the same as the privacy level of model $f_{\theta}(\bm{x};\theta)$ parameterised by $\Omega_{\theta}$ with prior $\tilde{\pi}_{0}(\theta)$,
@@
Recall that $f_{\psi}(\bm{x};\psi)$ and $\pi_{0}(\psi)$ are the likelihood and the prior with respect to the parameter space $\Omega_{\psi}$, and
$f_{\theta}(\bm{x};\theta)=f_{\psi}(\bm{x};t^{-1}(\theta))\cdot|J_{t^{-1}}(\theta)|,\qquad\tilde{\pi}_{0}(\theta)=\pi_{0}(t^{-1}(\psi))\cdot|J_{t^{-1}}(\theta)|,$
where $J_{t^{-1}}(\theta)$ is the Jacobian of the inverse function. Note that $f_{\theta}$ and $\tilde{\pi}_{0}$ are derived from $f_{\psi}$, $\pi_{0}$ through a change of measure under mapping $t$.
@@
$\implies$ $\,\mathbb{P}_{\pi_{\psi}}(t^{-1}(S)|\bm{X})\leq e^{\epsilon}\mathbb{P}_{\pi_{\psi}}(t^{-1}(S)|\bm{Z})+\delta,$
and vice versa, since the integral is preserved through the change of measure. Therefore, a pair of $(\epsilon,\delta)$ holds for a model parameterised in $\Omega_{\psi}$ if and only if it holds for that model parameterised in $\Omega_{\theta}$ through the diffeomorphism $t$. ∎
@@
where $\mu$ is the base measure for the measure space $\Omega_{\psi}$. That is, $\psi_{1}$ and $\psi_{2}$ lead to two likelihood functions that are equal almost everywhere. Since equal a.e. could create an issue in the DP inequality as the worst case could land on the discontinuity, we consider $f(\bm{x};\psi)$ to be continuous in $\bm{x}$. In which case, $\psi_{1}\sim\psi_{2}$ if and only if they represent the same likelihood function on the whole of $\mathcal{X}$.
@@
$\begin{array}[]{cccc}t:&\Omega_{\psi}&\rightarrow&\Omega_{\theta},\\ &\psi&\mapsto&[\psi].\end{array}$
If there is a way to choose every representative $\psi]$ such that $t$ is measurable, then we have the following proposition.
@@
Let $(\Omega_{\psi},\Sigma_{\psi},\mathtt{d}\psi)$ be a measure space such that the model with likelihood function $f_{\psi}(x;\psi)$ and prior $\pi_{0}(\psi)$ is non-identifiable. Suppose that the quotient map $t:\Omega_{\psi}\rightarrow\Omega_{\theta}\subset\Omega_{\psi}$ that reparameterises the model onto the measure space $(\Omega_{\theta},\Sigma_{\theta},\mathtt{d}\theta)$ is measurable, then the model parameterised in $\Omega_{\theta}$ has the same level of differential privacy as the model parameterised in $\Omega_{\psi}$.
@@
under the same construction of the quotient map $t$, pushforward measure $t_{*}(\mathbb{P}_{\pi_{0}})$ and its density $\tilde{\pi}_{0}$ with respect to $\mathtt{d}\theta$ as in the previous proof. Since the pushforward measure and the design of the quotient ensure that any integrals are preserved, the $\implies$ direction is trivial.
@@
$\frac{\pi_{\psi}(\psi|\bm{X})}{\pi_{\psi}(\psi|\bm{Z})}<e^{\epsilon},\quad\forall\psi\in B,$
@@
Without loss of generality, it is now sufficient to check that $(\epsilon,\delta)$ holds for all sets $A\in\Sigma_{\psi}$ such that
@@
For any set $A\in\Sigma_{\psi}$ such that (39) holds, let $S=t(A)$, note that any $\psi\in t^{-1}(S)\supseteq S$ also satisfies (39),
@@
Appendix F Extra Figures
Figure 3 is a replot of Figure 2 in the main text, where now the $\epsilon$ is plotted in log scale, and it is easier to see the polynomial decay in $\epsilon$ with respect to $n$.
Figure 3: Figure 2 plotted in log-scale. Plots of estimated $\epsilon_{n}$ under the setting of Linear Regression, Logistic Regression and Cauchy Regression as the dataset size $n$ varies. The pale lines without nodes are extrapolated estimations.
{acks}
[Acknowledgments] We would like to thank all the reviewers and the associate editor for reviewing and providing constructive comments! {funding} All authors were supported by the EPSRC research grant ”Pooling INference and COmbining Distributions Exactly: A Bayesian approach (PINCODE)”, reference (EP/X028100/1, EP/X028119/1, EP/X028712/1, EP/X027872/1).
LA, HD, MP and GOR were also supported by the UKRI grant, ”On intelligenCE And Networks (OCEAN)”, reference (EP/Y014650/1).